# Efficient Visual Transformer by Learnable Token Merging

### v1

@@ 5.3 Visualization Results
To study the effectiveness of LTM-Transformer in selecting informative tokens during the token merging process, we visualize the token merging masks in the first LTM-Transformer block of LTM-MobileViT-S for selected images from ImageNet in Figure 5. Each image is divided into $16\times 16$ tokens. For each example, we select only the most representative merged token that encapsulates the critical features of the objects in the image, and the merged token is a weighted average of several self-attention output tokens with the aggregation weights in the token merging mask. The input images are illustrated in the first row, and the heatmaps that visualize the aggregation weights in the token merging mask for the selected merged token are shown in the second row. The class labels for each image are presented at the bottom of each column. The results illustrate that the mask module in the LTM-Transformer block usually assigns higher aggregation weights to tokens covering the most representative and distinctive parts of the objects, which are often the most informative for classifying the images. In the example of the dhole in the first column, the LTM-Transformer block puts larger weights on the eyes and nose of the dhole. In the example of the hartebeest in the second column, the LTM-Transformer block puts larger weights on the twisted horns of the hartebeest. In the example of the racing car in the third column, the LTM-Transformer block puts larger weights on the wheel of the car. In the example of the Stethoscope in the fourth column, the LTM-Transformer block puts larger weights on the diaphragm of the stethoscope. These observations demonstrate that more informative tokens contribute more to the merged tokens with larger aggregation weights in the token merging process of the LTM-Transformer block.
6 Proofs
6.1 Proof of Proposition 3.2
Proof.
We first compute the gradient of $\phi(\tilde{X}_{i}(G),a^{\prime})$ with respect to $\tilde{X}_{i}(G)$ by $\nabla_{\tilde{X}_{i}(G)}\phi(\tilde{X}_{i}(G),a^{\prime})=\frac{2S_{i{a^{\prime}}}\sum\limits_{a=1}^{C}S_{ia}\left(\tilde{X}_{i}(G)-\tilde{\cal C}_{a}\right)-2S_{i{a^{\prime}}}\left(\tilde{X}_{i}(G)-\tilde{\cal C}_{a^{\prime}}\right)\sum\limits_{a=1}^{C}S_{ia}}{\left(\sum\limits_{a=1}^{C}S_{ia}\right)^{2}}.$
Using the definitions of $\gamma_{i}$ and $\zeta_{i}$ as $\gamma_{i}\coloneqq\sum\limits_{a=1}^{C}S_{ia}$ and $\zeta_{i}\coloneqq\sum\limits_{a=1}^{C}S_{ia}\tilde{\cal C}_{a}$ for $i\in[n]$, we have
$\nabla_{\tilde{X}_{i}(G)}\phi(\tilde{X}_{i}(G),a^{\prime})=\frac{2S_{i{a^{\prime}}}}{\gamma_{i}^{2}}\left(\gamma_{i}\tilde{\cal C}_{a^{\prime}}-\zeta_{i}\right).$
As a result, the gradient of $\textup{IBB}(G)$ with respect to $G$ is computed as follows: $\nabla_{G}\textup{IBB}(G)$ $=\frac{1}{n}\sum\limits_{i=1}^{n}\sum\limits_{a=1}^{C}\sum\limits_{b=1}^{C}Z_{i}\nabla_{\tilde{X}_{i}}\phi(\tilde{X}_{i},a)\phi(X_{i},b)\log{\phi(X_{i},b)}$ $-\frac{1}{n}\sum\limits_{i=1}^{n}\sum\limits_{a=1}^{C}\sum\limits_{y=1}^{C}Z_{i}\nabla_{\tilde{X}_{i}}\phi(\tilde{X}_{i},a){\rm 1}\kern-2.25pt{\rm I}_{\left\{y_{i}=y\right\}}\log{Q(\tilde{X}\in a|Y=y)}$ $=\frac{1}{n}\sum\limits_{i=1}^{n}\sum\limits_{a=1}^{C}Z_{i}\nabla_{\tilde{X}_{i}}\phi(\tilde{X}_{i},a)\psi_{i,a}$ $=\frac{2}{n}\sum\limits_{i=1}^{n}\sum\limits_{a=1}^{C}Z_{i}\frac{S_{i{a}}}{\gamma_{i}^{2}}\left(\gamma_{i}\tilde{\cal C}_{a}-\zeta_{i}\right)\psi_{i,a},$ where $\psi_{i,a}\coloneqq\sum\limits_{b=1}^{C}\phi(X_{i},b)\log{\phi(X_{i},b)}-\sum\limits_{y=1}^{C}{\rm 1}\kern-2.5pt{\rm I}_{\left\{y_{i}=y\right\}}\log{Q(\tilde{X}\in a|Y=y)}$.
6.2 Proof of Theorem 3.1
We need the following two lemmas before the proof of Theorem 3.1. It is noted that we abbreviate $\tilde{X}(G)$ and $\tilde{X}_{i}(G)$ as $\tilde{X}$ and $\tilde{X}_{i}$ in the sequel.
Lemma 6.1.
$I(\tilde{X},X)$ $\leq\frac{1}{n}\sum\limits_{i=1}^{n}\sum\limits_{a=1}^{C}\sum\limits_{b=1}^{C}\phi(\tilde{X}_{i},a)\phi(X_{i},b)\log{\phi(X_{i},b)}$
$-\frac{1}{n^{2}}\sum\limits_{i=1}^{n}\sum\limits_{j=1}^{n}\sum\limits_{b=1}^{C}\phi(X_{i},b)\log{\phi(X_{j},b)}$ (4)
Lemma 6.2.
$I(\tilde{X},Y)\geq\frac{1}{n}\sum\limits_{a=1}^{C}\sum\limits_{y=1}^{C}\sum\limits_{i=1}^{n}\phi(\tilde{X}_{i},a){\rm 1}\kern-2.5pt{\rm I}_{\left\{y_{i}=y\right\}}\log{Q(\tilde{X}\in a|Y=y)}$ (5)
Proof of Theorem 3.1.
We note that $\textup{IB}({\cal W})=I(\tilde{X},X)-I(\tilde{X},Y)$. Then $\textup{IB}({\cal W})\leq\textup{IBB}({\cal W})-C_{0}$ follows by the upper bound for $I(\tilde{X},X)$ in Lemma 6.1 and the lower bound for $I(\tilde{X},Y)$ in Lemma 6.2. Here $C_{0}=\frac{1}{n^{2}}\sum\limits_{i=1}^{n}\sum\limits_{j=1}^{n}\sum\limits_{b=1}^{C}\phi(X_{i},b)\log{\phi(X_{j},b)}$. ∎
Proof of Lemma 6.1.
By the log sum inequality, we have
$I(\tilde{X},X)$
$=\sum\limits_{a=1}^{C}\sum\limits_{b=1}^{C}\textup{Pr}\left[\tilde{X}\in a,X\in b\right]\log{\frac{\textup{Pr}\left[\tilde{X}\in a,X\in b\right]}{\textup{Pr}\left[\tilde{X}\in a\right]\textup{Pr}\left[X\in b\right]}}$
$\leq\frac{1}{n^{2}}\sum\limits_{i=1}^{n}\sum\limits_{j=1}^{n}\sum\limits_{a=1}^{C}\sum\limits_{b=1}^{C}\phi(\tilde{X}_{i},a)\phi(X_{i},b)\left(\log\left(\phi(\tilde{X}_{i},a)\phi(X_{i},b)\right)\right.$
$\left.-\log\left(\phi(\tilde{X}_{i},a)\phi(X_{j},b)\right)\right)$
$=\frac{1}{n^{2}}\sum\limits_{i=1}^{n}\sum\limits_{j=1}^{n}\sum\limits_{a=1}^{C}\sum\limits_{b=1}^{C}\phi(\tilde{X}_{i},a)\phi(X_{i},b)\log{\phi(X_{i},b)}$
$\phantom{=}-\frac{1}{n^{2}}\sum\limits_{i=1}^{n}\sum\limits_{j=1}^{n}\sum\limits_{a=1}^{C}\sum\limits_{b=1}^{C}\phi(\tilde{X}_{i},a)\phi(X_{i},b)\log{\phi(X_{j},b)}$
$=\frac{1}{n}\sum\limits_{i=1}^{n}\sum\limits_{a=1}^{C}\sum\limits_{b=1}^{C}\phi(\tilde{X}_{i},a)\phi(X_{i},b)\log{\phi(X_{i},b)}$
$\phantom{=}-\frac{1}{n^{2}}\sum\limits_{i=1}^{n}\sum\limits_{j=1}^{n}\sum\limits_{a=1}^{C}\sum\limits_{b=1}^{C}\phi(\tilde{X}_{i},a)\phi(X_{i},b)\log{\phi(X_{j},b)}.$ (6)
Proof of Lemma 6.2.
Let $Q(\tilde{X}|Y)$ be a variational distribution. We have
$I(\tilde{X},Y)$
$=\sum\limits_{a=1}^{C}\sum\limits_{y=1}^{C}\textup{Pr}\left[\tilde{X}\in a,Y=y\right]\log{\frac{\textup{Pr}\left[\tilde{X}\in a,Y=y\right]}{\textup{Pr}\left[\tilde{X}\in a\right]\textup{Pr}\left[Y=y\right]}}$
$=\sum\limits_{a=1}^{C}\sum\limits_{y=1}^{C}\textup{Pr}\left[\tilde{X}\in a,Y=y\right]\log{\frac{\textup{Pr}\left[\tilde{X}\in a|Y=y\right]Q(\tilde{X}\in a|Y=y)}{\textup{Pr}\left[\tilde{X}\in a\right]Q(\tilde{X}\in a|Y=y)}}$
$\geq\sum\limits_{a=1}^{C}\sum\limits_{y=1}^{C}\textup{Pr}\left[\tilde{X}\in a,Y=y\right]\log{\frac{\textup{Pr}\left[\tilde{X}\in a|Y=y\right]}{Q(\tilde{X}\in a|Y=y)}}$
$+\sum\limits_{a=1}^{C}\sum\limits_{y=1}^{C}\textup{Pr}\left[\tilde{X}\in a,Y=y\right]\log{\frac{Q(\tilde{X}\in a|Y=y)}{\textup{Pr}\left[\tilde{X}\in a\right]}}$
$=\textup{KL}\left(P(\tilde{X}|Y)\middle\|Q(\tilde{X}|Y)\right)$
$+\sum\limits_{a=1}^{C}\sum\limits_{y=1}^{C}\textup{Pr}\left[\tilde{X}\in a,Y=y\right]\log{\frac{Q(\tilde{X}\in a|Y=y)}{\textup{Pr}\left[\tilde{X}\in a\right]}}$
$\geq\sum\limits_{a=1}^{C}\sum\limits_{y=1}^{C}\textup{Pr}\left[\tilde{X}\in a,Y=y\right]\log{\frac{Q(\tilde{X}\in a|Y=y)}{\textup{Pr}\left[\tilde{X}\in a\right]}}$
$=\sum\limits_{a=1}^{C}\sum\limits_{y=1}^{C}\textup{Pr}\left[\tilde{X}\in a,Y=y\right]\log{Q(\tilde{X}\in a|Y=y)}+H\left(P(\tilde{X})\right)$
$\geq\sum\limits_{a=1}^{C}\sum\limits_{y=1}^{C}\textup{Pr}\left[\tilde{X}\in a,Y=y\right]\log{Q(\tilde{X}\in a|Y=y)}$
$\geq\frac{1}{n}\sum\limits_{a=1}^{C}\sum\limits_{y=1}^{C}\sum\limits_{i=1}^{n}\phi(\tilde{X}_{i},a){\rm 1}\kern-2.5pt{\rm I}_{\left\{y_{i}=y\right\}}\log{Q(\tilde{X}\in a|Y=y)}.$ (7)
6.3 Computation of $Q^{(t)}(\tilde{}\mathbf{X}|Y)$
The variational distribution $Q^{(t)}(\tilde{}\mathbf{X}|Y)$ can be computed by
$Q^{(t)}(\tilde{X}\in a|Y=y)$ $=\textup{Pr}\left[\tilde{X}\in a|Y=y\right]$
$=\frac{\sum\limits_{i=1}^{n}\phi(\tilde{X}_{i},a){\rm 1}\kern-2.5pt{\rm I}_{\left\{y_{i}=y\right\}}}{\sum\limits_{i=1}^{n}{\rm 1}\kern-2.5pt{\rm I}_{\left\{y_{i}=y\right\}}}.$ (8)
@@ 7 Conclusion
In this paper, we propose a novel transformer block, Transformer with Learnable Token Merging, or LTM-Transformer. LTM-Transformer blocks perform token merging so as to render a transformer network with less FLOPs and faster inference speed. A LTM-Transformer block generates an informative token merging mask for token merging in a learnable manner, which is inspired by the reduction of the Information-Bottleneck loss. A network with LTM-Transformer blocks can be trained with standard SGD, and it enjoys a reduction of IB loss and reduced FLOPs while maintaining a compelling prediction accuracy. We demonstrate the effectiveness of LTM-Transformer by replacing all the transformer blocks in several popular visual transformers with LTM-Transformer blocks. Extensive experiments on various tasks demonstrate the effectiveness of LTM-Transformer.

### v2

@@ 4.4.5 Visualization Results
To study the effectiveness of LTM-Transformer in selecting informative tokens during the token merging process, we visualize the token merging masks in the first transformer block of LTMP-MobileViT-S and LTM-MobileViT-S for particular target tokens in selected images from ImageNet in Figure 4. Each image is divided into $16\times 16$ tokens. For each example, we select only the most representative target token that encapsulates the critical features of the objects in the image, and the target token is a weighted average of several self-attention output tokens with the token merging weights in the token merging mask. The input images are illustrated in the first row. Every target token is marked in red boxes in the first row, which is calculated by $\tilde{X}(G^{(\ell)})=\left(Z^{\top}G^{(\ell)}\right)^{\top}$ as described in Section 3.1. The third row illustrates the token merging mask $G^{(\ell)}$ generated by the LTM for the target tokens with heatmaps. We remark that the token merging masks for LTM and LTMP contain the token merging weights for these two methods. The token merging masks generated by LTMP for the same target tokens are illustrated in the second row. The class labels for each image are presented at the bottom of each column. Tokens that are more semantically similar to the target token are expected to receive larger token merging weights in the token merging masks, so that the original tokens can contribute to the target token in accordance to their informativeness with respect to the target token. The visualization results illustrate that the mask module in the LTM-Transformer block usually assigns higher merging weights to tokens that are more semantically similar to the target token. In contrast, LTMP assigns uniform merging weights to all the selected tokens, which usually contain some undesirable tokens that are not semantically similar or even relevant to the target tokens. In the example of the dhole in the first column, the LTM-Transformer block assigns larger merging weights to tokens around the eyes of the hole as the target token is also around the eyes. In contrast, the competing baseline LTMP mistakenly merges tokens on the ears of the hole into the target token. In the example of the hartebeest in the second column, the LTM-Transformer block assigns larger weights to tokens on the twisted horns of the hartebeest as the target token is also on the twisted horns of the hartebeest. The uneven distribution of token merging weights computed by LTM is consistent with the goal of informative token merging in our LTM-Transformer, where more informative original tokens, whose features are more similar to that of the target token, should receive higher token merging weights than other original tokens. In contrast, the competing baseline LTMP mistakenly merges tokens in the background around the twisted horns of the hartebeest. In the example of the racing car in the third column, the LTM-Transformer block assigns larger weights to tokens on the wheel of the car as the target token is also on the wheel. In contrast, the competing baseline LTMP mistakenly merges tokens around the light of the car. In the example of the Stethoscope in the fourth column, the LTM-Transformer block assigns larger weights to tokens on the diaphragm of the stethoscope. In contrast, the competing baseline LTMP mistakenly merges tokens on the top of the diaphragm of the stethoscope in the background.
In addition, the visualization results illustrate that LTMP usually assigns larger weights to tokens that are more semantically similar to the target tokens. For instance, in the example of the hartebeest, tokens on the left horn of the hartebeest have larger token merging weights than tokens on the right horn of the hartebeest because the target token is on the left horn, which exhibits visual patterns that are not shared on the right horn. Similarly, in the example of the stethoscope, tokens around the center of the diaphragm have larger merging weights than tokens on the boundary of the diaphragm because the target token is in the center of the diaphragm, which exhibits visual patterns that are not shared on the boundary of the diaphragm.
4.4.6 Study on the Impacts of Compression Ratio at Different Layers of the LTM-Transformer
In our experiments, the compression ratio $r$ for all the layers in the LTM-Transformer is set to a fixed value of $0.7$. To study how the compression ratio at different layers in the LTM-Transformer influences the IB loss at the corresponding layer, we train ablation models by changing the compression ratios at different layers from $0.7$ to $0.3$, $0.5$, and $0.9$. For each of the ablation models, we only change the compression ratio at a specific layer and keep the compression ratios for other layers fixed at $0.7$. The ablation study is performed for all $9$ layers of the LTM-MobileViT-S in the fine-tuning setup, and all ablation models are fine-tuned for $50$ epochs. Let $\ell$ denote the layer index where the compression ratio is changed. The IB loss at the corresponding layer where the compression ratio is changed and the top-1 accuracy of the ablation models are shown in Table VIII. It is observed that reducing the compression ratio from $r=0.9$ to $r=0.7$ can reduce the IB loss at the corresponding layer and increase the top-1 accuracy, benefiting from merging redundant tokens. However, further reducing the compression ratio to $r=0.5$ and $r=0.3$ leads to an increased IB loss and reduced top-1 accuracy, especially at bottom layers where $\ell\leq 5$. This is attributed to the excessive merging of tokens, potentially resulting in the loss of informative features that are critical for the classification task.
Layer $r=0.3$ $r=0.5$ $r=0.7$ $r=0.9$
Index $\ell$ Top-1 IB Loss Top-1 IB Loss Top-1 IB Loss Top-1 IB Loss
1 78.22 0.00214 78.70 0.00190 79.05 0.00178 79.02 0.00185
2 78.43 -0.00137 78.85 -0.00140 79.05 -0.00156 79.01 -0.00148
3 78.55 -0.00622 78.90 -0.00630 79.05 -0.00648 78.98 -0.00641
4 78.57 -0.00725 79.05 -0.00742 79.02 -0.00762 78.98 -0.00750
5 78.75 -0.00938 78.83 -0.00950 79.05 -0.01005 79.00 -0.00977
6 78.95 -0.01185 79.02 -0.01197 79.05 -0.01204 79.04 -0.01199
7 79.00 -0.01430 79.04 -0.01464 79.05 -0.01488 79.01 -0.01486
8 79.00 -0.01525 79.03 -0.01556 79.05 -0.01592 79.05 -0.01592
9 79.04 -0.01718 79.03 -0.01704 79.05 -0.01725 79.00 -0.01728
TABLE VIII: Ablation study on the impacts of compression ratio in reducing the IB loss at different layers of the LTM-MobileViT-S. All ablation models are fine-tuned for 50 epochs. $\ell$ denotes the layer index where the compression ratio is changed. The IB loss at the $\ell$-th layer and the top-1 accuracy of the ablation models are shown.
In addition, we have performed an ablation study on evaluating the transfer learning capability of the LTM-Transformer in Section 2.1 of the supplementary. The results show that LTM does not affect the transfer learning capability of the original model. Moreover, we have shown that the parameter-efficient post-training method, LoRA [48], can significantly reduce the training costs of the LTM-Transformer without scarfing the performance in Section 2.2 of the supplementary. In Section 2.3 of the supplementary, we have shown that the LTM-Transformer exhibits faster inference speed than the models compressed by competing token merging methods across different batch sizes. In Section 3 of the supplementary, we have discussed the potential future work of this paper for improved training efficiency.
@@ 5 Conclusion
In this paper, we propose a novel transformer block, Transformer with Learnable Token Merging, or LTM-Transformer. LTM-Transformer blocks perform token merging so as to render a transformer network with less FLOPs and faster inference speed. A LTM-Transformer block generates an informative token merging mask for token merging in a learnable manner, which is inspired by the reduction of the Information-Bottleneck loss. A network with LTM-Transformer blocks can either be trained from scratch or fine-tuned from the pre-trained backbone, and it enjoys a reduction of IB loss and reduced FLOPs while maintaining a compelling prediction accuracy. We demonstrate the effectiveness of LTM-Transformer by replacing all the transformer blocks in several popular visual transformers with LTM-Transformer blocks. Extensive experiments on various tasks demonstrate the effectiveness of the LTM-Transformer.